Three years ago Thokozile borrowed R7 500 from Alfred. The condition was that she would pay him back in seven years’ time at an interest rate of 11,21% per year, compounded semi-annually. Six months ago she also borrowed R25 000 from Alfred at 9,45% per year, compounded monthly. Thokozile would like to pay off her debt four years from now .

After seeing what she must pay Alfred, Thokozile decides to reschedule her debt as two equal payments: one payment now and one three years from now. Alfred agrees on condition that the new agreement, that will run from now, will be subjected to 10,67% interest, compounded quarterly. The amount that Thokozile will pay Alfred three years from now is

[1] R22 286,88.

[2] R25 103,93.

[3] R32 500,00.

[4] R21 171,35.

[5] none of the above. 


1
Expert's answer
2021-05-28T09:49:27-0400

Present value of debt:

"PV=FV\/(1+r)^n"

where FV is future value,

r is interest rate,

n is number of periods in the future.


"PV_1=\frac{7500}{(1+0.1121/2)^{4\cdot2}}=R\ 4848.25"


"PV_2=\frac{25000}{(1+0.0945/12)^{3.5\cdot12}}=R\ 17982.90"


Payment now (payment of 1st debt for three years and 2nd debt for six months):

"PV_1+PV_2=4848.25+17982.90=R\ 22831.15"

The rest of debt:

"25000+7500-22831.15=R\ 9668.85"


Payment three years from now:

"9668.85(1+0.1067)^{3\cdot4}=R\ 32500"


Answer: [3] R32 500,00